a beta particle of energy 1MeV enters a thin window detector tube filled with methane and organ and initiates an avalanche of 5×10^10 Ion pairs the average energy lose per ion pair creation is 50 eV. the multiplication factor is
1
Expert's answer
2019-04-30T09:52:59-0400

the multiplication factor is equal to


"\frac {50 \times 10^{10} \times 50} {1\times 10^6} =2.5 \times 10^7"


Answer:

"2.5 \times 10^7"